OpenStudy (anonymous):

Find the limit or verify that the limit does not exist: lim(x,y)--->(0,2) f(x,y)=(x+y-2)/(x^2+y^2+2xy-4)

OpenStudy (anonymous):

\[\lim_{(x,y)\to(0,2)}\frac{x+y-2}{x^2+y^2+2xy-4}\] First thing that stands out to me is that \[x^2+2xy+y^2=(x+y)^2\] so you have \[\lim_{(x,y)\to(0,2)}\frac{(x+y)-2}{(x+y)^2-4}\] i.e. a difference of squares in the denominator, which can be factored: \[\lim_{(x,y)\to(0,2)}\frac{(x+y)-2}{((x+y)-2)((x+y)+2)}\\ \lim_{(x,y)\to(0,2)}\frac{1}{x+y+2}\] No discontinuities/indeterminate forms to worry about, so you can evaluate directly.